OpenStudy (anonymous):

Hi would appreciate some assistance with this. I am supposed to simplify and find an extraneous solution (if there is one). (7a^2 b^3)^1/2 (7a^2b^3)^-1/2

OpenStudy (preetha):

Irish, can you explain...

OpenStudy (anonymous):

Is that multiplication between the two expressions? @dawidskalkowski

OpenStudy (anonymous):

I just did the math, and I also got the answer 1. I'm just double-chcking. And yes genius, it is.

mathslover (mathslover):

\((7a^2 b^3)^\frac{1}{2} (7a^2b^3)^{\frac{-1}{2}} \) Note : \(a^{-m} = \cfrac{1}{a^m} \) So, \(\large { (7a^2 b^3 ) ^\frac{-1}{2} = \cfrac{1}{(7a^2 b^3 ) ^\frac{1}{2}} }\) Now, we get : \(\large {(7a^2 b^3)^\frac{1}{2} (7a^2b^3)^{\frac{-1}{2}} \implies (7a^2 b^3 ) ^\frac{1}{2} \times \cfrac{1}{(7a^2 b^3 ) ^\frac{1}{2}} } \) \(\large{\cancel{(7a^2 b^3) ^\frac{1}{2}} \times \cfrac{1}{\cancel{(7a^2 b^3 ) ^\frac{1}{2}}} }\) = \(\color{blue}{\large{1}} \)

OpenStudy (irishboy123):

that looks good basically X ^(1/2) * X ^ (-1/2) = X ^ 0 = 1

mathslover (mathslover):

Right, it can also be explained further in the form of identity : \(\large a^m \times a^n = a^{m+n}\) If m = -n It becomes : \(\large a^{\color{blue}{-n}} \times a^{n} = a^{\color{blue}{-n } + n }= a^0 = 1\)